Reports of 51595
AdV-SGD (FDS commissioning)
vardaro, bawaj - 17:22 Monday 08 August 2022 (56689) Print this report
Comment to SQZ_FLT: bugfix and preparation for a long test run (56656)

After the restore of yesterday the Automation worked well up to 6:15 UTC of this morning. Here the cavity unlocked and was not more able to relock again (fig1) . We did the following observation:

  • after the unlock the system fall in down and discharged the correction of all the mirrirs (FCIM, FCEM GR_M5 and GR_M7) but some kicks to the mirror occured. After that the transmission was at 4V maximum (fig2)
  • the threshold for the lock was set at 6.2 V thus the system did not relock on green anymore
  • sometimes after the lock on green we noticed a lot of frequency noise and the BPC2 based on dither lines on FCIM and FCEM did not work well. 

To solve it we are implementing the following solutions:

  • increase the SQZ_CC_Ctrl GAIN from 0.12 to 0.5 to reduce frequency noise (temorary bug fixing)
  • lower the threshold for engage the lock on green from 6.2 to 5 i.e. the 75% of the maximum transmission
  • insert in the green+IR locking a state to re-align the green if the transmission is too low, because in that condition it is difficult to find the IR
  • try to discharge the mirrors correction without sending kicks to the actuators. A possible solution is tested (fig 3).

We are impementing these upgrades in the SQZ_FLT.py code and tomorrow we will test them. 

Note that other unlocks were triggered by other issues in the automation. These kind of unlocks was less critical thus we started from the bigger problem

Images attached to this comment
AdV-INJ (Input Mode Cleaner cavity)
melo - 16:44 Monday 08 August 2022 (56688) Print this report
Comment to INJ throughput losses (06 may to 28 june) (56303)

There is a mistake in the text of this entry, the top graph of the first attached figure shows the power towards the IMC cavity, not the transmission of the Mode Cleaner.

AdV-INJ (Input Mode Cleaner cavity)
melo, chiummo - 16:34 Monday 08 August 2022 (56687) Print this report
MC throughput losses before and after CHRoCC installation

We analyzed the throughput losses of the Mode Cleaner before and after the CHRoCC installation in PR and SR towers.

- BEFORE CHRoCC INSTALLATION

The 1st and 2nd attached figures show the trend for the period of May 6th 2022 to July 16th 2022.

In the first figure we can see three graphs: the top graphs shows the power in the output of EIB, that is, the power in the input of the IMC. The middle graph shows the PSL to INJ throughput, i.e.m the ratio between the calculated power at the output of the PMC (we have estimated 7% losses from the output of the PMC to the output of EIB) and the power at the input of the interferometer (which is estimated to be 7.5% less w.r.t the output of the IMC due to SIB1 losses), and bottom graph shows the throughput losses of the Mode Cleaner cavity, i.e., the intracavity losses, which is around 17% for the time window analyzed.

The second figure, on the other hand, shows the mismatch calculated for the same period. The mean value of mismatch is around 8.5%.

- AFTER CHRoCC INSTALLATION

The same plots have been produced for the period of July 23rd 2022 to August 8th 2022, i.e., after the two CHRoCCs installation.

The third plot shows the EIB output power (top graph), PSL to INJ throughput (middle graph) and the IMC throughput losses (bottom graph). Here we can estimate that the throughput losses are around 15.5%, which is 1.5% less w.r.t the value before the CHRoCC installation.

The fourth plot shows the mismatch for this period. As we can see, the mismatch has reached a mean value of around 9.5%, representing an increase of 1% w.r.t the previous value.

 

Images attached to this report
Detector Operation (Operations Report)
gherardini - 16:00 Monday 08 August 2022 (56681) Print this report
Operator Report - Morning shift

This morning the scheduled activity of study of angular d.o.f. on CITF was carried out by Diego and Manuel, the work went on the whole shift without any major problem, this activity is still ongoing...

Sub-system reports

DET
DET_MAIN switched state to FAST_SHUTTER_DEAD at around 4:28UTC, I restored it back to SHUTTER_CLOSED.

Other
yesterday at 15:08UTC the SQB2 position loop was opened by the guardian, controls properly closed.

Images attached to this report
AdV-ISC (Commissioning up to first full interferometer lock)
ruggi - 15:41 Monday 08 August 2022 (56685) Print this report
Alignment figures of merit

This week a few ISC shifts have been allocated to the optimization of alignment in CITF. For the evaluation of alignment quality, a few figures of merit have been identified.

In order to see the 'natural' variation of some figure of merit, two gps have been taken into account:

1343811518  Sat Aug  6 08:58:20 2022

1343980618  Mon Aug  8 07:56:40 2022

As far as I understand, all the settings should be the same for the two locks, but I could have lost something.

According to fig 1, the lock of Saturday is a bit better: B4_DC and B4_12MHz_mag are higher, B4_112MHz_mag is almost the same, B1p_DC is lower. The information coming from B4 phase camera confirms the impression: the overlap between carrier and sidebands (fig 2) is higher.

A possible variation in the quality of the alignment is monitored by the demodulation of B7_DC and B8_DC at the angular dither lines on ITMs. The signals are shown in fig 3; they refer to pre-alignments occurred 5000 s before the time of the CITF lock. In both the cases, the signals are not too far from zero, but one is better than the other. That is the one of Monday, corresponding to the worse CITF.

This analysis goes together with a much more evident observation: in the last days the figures of merit taken into account here have been dramatically influenced by a fine tuning of TCS actuators. Any undesired variation of those figures of merit, observed from time to time, could be attributed to some undesired variation of the parameters on which the TCS actuators have an impact. For this reason, we have to be careful in using of the same figures of merit for evaluating the quality of the alignment.

 

Images attached to this report
AdV-TCS (CO2 laser projector)
nardecchia - 15:00 Monday 08 August 2022 (56686) Print this report
Comment to NI CH center position check with NI RH and HWS DET (56668)

I have to correct the sentence:

''Thus, NI CH and RH are centered within 12*12.28/1024*11~2 mm in X and within  33*12.28/1024~4.5 mm in Y.''

in

''Thus, NI CH and RH are centered within 12*12.28/1024*13~2 mm in X and within  33*12.28/1024*13~5 mm in Y.''

I have mistakenly used the HWS-INJ magnification (11) instead of the HWS-DET one (13).

AdV-DET (Commissioning)
gouaty, hui, mwas - 9:55 Monday 08 August 2022 (56684) Print this report
Comment to OMC mode matching and benches alignment checks (56654)

During the activity of friday morning (OMC mode matching), the bench angular setpoints have evolved from 136 to 29 urad in TX , and from 25 to 32 urad in TY. The change in TY seems rather unsignificant, but the change in TX is quite substantial. We see that at the same time the beam spot is moving on the B1p camera. The main effect is a horizontal motion on the camera by 300 um.

Considering the size of the photodiodes (3 mm aperture), a motion of 300 um should not be an issue, but it may still be worth to check the centering of the beam on the photodiodes and the quadrants on SDB2.

Also probably the alignment of the B1s QPD on EDB will have to be checked.

Images attached to this comment
AdV-ISC (Commissioning up to first full interferometer lock)
jerome.degallaix - 9:14 Monday 08 August 2022 (56683) Print this report
Comment to ISC/TCS shift: Chrocc power tuning in CITF 1f (56672)

As noticed in the post, the CHRoCC voltage of 0.43V is a really good starting point. As displayed in the VIM pages, the SB power recycling gain for the 6 MHz was ~65 and ~40 for the 56 MHz. It has to be compared to 40 and 20 without the CHRoCC (see VIR-0369A-22 for example)

Images attached to this comment
AdV-DAQ (Data Acquisition and Global Control)
masserot - 8:36 Monday 08 August 2022 (56682) Print this report
Comment to TolmFrambeBuilde RT-FIFO readout frequency - Tpro_elapsed_time side effet (56662)

All the TolmFrameBuilder servers building a frame at 5Hz are now running with  the DAQ_FIFO_DUMP_FREQ    set to 400Hz.

The operations were performed between 2022-08-08-04h26m-UTC and 2022-08-08-04h33m-UTC

Detector Operation (Operations Report)
gherardini - 8:04 Monday 08 August 2022 (56680) Print this report
Operator Report - Initial Status
ITF State: cavities locked on infrared.
Quick Summary: all suspension loops closed; all SBEs loops closed with the exception of SQB2 position control.
No activity ongoing.
Images attached to this report
AdV-SGD (FDS commissioning)
vardaro - 0:28 Monday 08 August 2022 (56679) Print this report
Comment to SQZ_FLT: bugfix and preparation for a long test run (56656)

The automation worked well up to 14:48 UTC of 7th August, then the filter cavity unlocked and started to lock/unlock up to 15 UTC see Fig1. A detailed entry to this behaviour will follow/

After that we experienced an INJ unlock and the ITF ML set point moved by about 0.3V (see Fig2). After that the PLLs were no more able to lock thus I had to tune the SQZ laser set point to lock the PLL in this working point (new set points in Fig3). Then I realigned the Green beam and the IR beam to the filter cavity and I re engaged the automation at 21:45 UTC see Fig4.

Images attached to this comment
Radius of Curvature Correction (Internal CHRoCC)
bersanetti - 21:25 Sunday 07 August 2022 (56678) Print this report
PR CHRoCC setpoint back to 0.43 V

In preparation for tomorrow's activities, the PR CHRoCC setpoint was moved back to V_DAC = 0.43 V at 19:22:00 UTC.

AdV-TCS (CO2 laser projector)
nardecchia - 20:05 Sunday 07 August 2022 (56677) Print this report
Comment to NI CH center position check with NI RH and HWS DET (56668)

The check of the last movement has been done today afternoon.

With the ITF in down, the NI CH has been switched ON at high power and the minimum OPL coordinates have been compared with the maximum RH OPL ones (see the attached figure).

The results are summarized in the following table:

 

X [pixels]

Y [pixels]

NI RH

729+-6

459+-4

NI CH

727+-4

464+-24

Difference

2+-10

5+-28

Thus, NI CH and RH are centered within 12*12.28/1024*11~2 mm in X and within  33*12.28/1024~4.5 mm in Y.

The ITF has been left locked on IR.

 

Images attached to this comment
AdV-TCS (CO2 laser projector)
nardecchia, taranto - 18:02 Sunday 07 August 2022 (56676) Print this report
Check of the WI CO2 actuators alignment on the WI CP

The goal of this activity was to check the alignment of the WI CO2 actuators on the WI CP.

The ITF has been put in down, the HWS-INJ acquisition started and the WI CH has been switched ON at high power at  06.55 UTC (WF 21). The CH beam appeared misaligned wrt the RH in both directions as you can see in the first attached figure (see entry 56673 for the WI TM center identification through RH).

Thus, the steering mirror on the CO2 bench has been iteratively moved in different steps to recover the centering of the CO2 beam actuators. 

After different steering mirror movements performed during the morning, the final check has been done with a new HWS-INJ acquisition.

The WI CH has been switched ON again at high power at 12.23 UTC (WF 10). Since the curvature is well visible and the coordinates are stable, it was switched off at 12.34 UTC. To be sure, at 13.12 UTC we switched on again the WI RH @15 W (WF 102) (it was switched off at 14.55 UTC) and the final configuration on the WI is shown in figure 2.

In the following table, the today WI RH and CH coordinates on the WI and their difference are reported.

 

X [pixels]

Y [pixels]

X [mm]

Y [mm]

WI RH

617 +- 2

423 +- 4

13.8 +- 0.3

-11.7 +- 0.5

WI CH

610 +- 3

421 +- 13

12.9 +- 0.4

-12 +- 2

Difference

7 +- 5

2 +- 17

0.9 +- 0.7

0.3 +- 2.2

We can conclude that the WI CH is well aligned wrt the WI center.

Images attached to this report
Detector Operation (Operations Report)
menzione - 16:01 Sunday 07 August 2022 (56675) Print this report
Operator Report - Daily shift

Upon arrival I found ITF locked on IR. I easily locked the CITF (LOCKED_ARMS_BEAT_DRMI_1F) making a cross.
After the unlock (?) I tried again but the WEST ALS didn't lock anymore (no flashes on WE_Cam). Probably a reset will need at WE.
At 06:40 UTC the planned TCS activity on RH was carried out from remote by Nardecchia, Taranto started and went on without major problems for the whole shift. 
I unlocked and relocked on IR. During the relock I had to realign NI and WI.

 ITF left unlocked.

No particular operations to report

Images attached to this report
Detector Operation (Operations Report)
menzione - 8:46 Sunday 07 August 2022 (56674) Print this report
Operator Report - Initial shift

ITF State: ITF Locked on IR.
ITF Mode: COMMISSIONING
Quick Summary: IMC and RFC locked; all Suspensions loops closed, all SBE loops closed.
Activities ongoing since this morning: No activity ongoing

No DMS Events to report

 

Sub-system reports

EnvMoni
Meteo station still down, red flags on DMS

Images attached to this report
AdV-TCS (Ring Heater)
lumaca, nardecchia - 22:46 Saturday 06 August 2022 (56673) Print this report
WI RH ON

After the activity described in 56672, the WI RH has been switched ON in order to check the centering of the WI CO2 actuators, activity planned for tomorrow.
The ITF has been put in down state, the HWS-INJ acquisition started and WI RH has been switched on @15W at 17.50 UTC and switched OFF at 19.44 UTC. The coordinates of the WI mirror center are summarized in the table below (see also attached figure):  

WI RH 

row [pixels]

col [pixels]

Y coord [mm]

X coord [mm]

 

429+-6

615+-4

1.23+-0.04

-0.99+-0.07


The ITF has been left locked on the IR.
 

Images attached to this report
AdV-ISC (Commissioning up to first full interferometer lock)
bersanetti, lumaca, pinto, menzione - 18:21 Saturday 06 August 2022 (56672) Print this report
ISC/TCS shift: Chrocc power tuning in CITF 1f

Task of the shift was to continue the Chrocc power tuning while locked on CITF.

At the beginning of the shift the ITF was locked by the operator, after the usual first pre-alignment, he managed to lock the CITF and, for a quick test, it was possible to go on with the subsequent steps of the acquisition up to the step after 480 mW of arms power, then it unlocked during the jump to CARM null 3f. No phase or gain tuning was performed, but some fine tuning will be eventually needed.

At 7.45 UTC we started the planned activity: with the last Chrocc power and DAS settings set by Diana we locked the CITF in 1f signals by putting a positive (+) CARM offset of +3000, in order to perform an arms scan in this working point.

After a minor adjustment of some LSC gain and MICH/SRCL demodulation phase, we started the scan:

8.06.35 UTC: 1st Scan CARM set +38.000 Hz with 60 sec ramp: the scan worked, but rigth after reaching the final offset value the CITF unlocked. So for the subsequent scans, we decided to use a lower offset value, i.e. +35000 Hz.

8.14.00 UTC: 2nd scan +35.000 Hz 60 sec ramp.
8.16.57 UTC: CITF unlocked on the way back to +3000 Hz.

8.47.05 UTC: 3rd scan +35.000 Hz 60 sec ramp. This time the CITF unlocked during the ramp to 35.000 Hz offset.

results of the first two scans are reported respectively in Fig.s 1 and 2.


N.B. in the last lock, while we were in the scan condition (positive offsets) we experienced a general improvement of the figures of merit that we monitor during the lock, e.g. B4 DC reached 2.61 mW; the 12 MHz SB 0.138 mW.
Also the phase cameras on B4 showed a more circular and symmetric shape, while the ones on B1p got quite worse (see Fig.s 3 and 4)


%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

After these scan tests, we relocked the CITF in the usual working conditions (negative CARM Offset -3000 Hz), in order to start the Chrocc power tuning activity.

Phase cameras image at this instance (9.00 UTC is reported in Fig.5)

At 9.04.16 UTC: Diana made the first movement setting the Chrocc Voltage from 0.43 to 0.36 V.

9.19.00 UTC: we noticed a general worsening trend, as the aforementioned figures of merit were decreasing and the fringe was getting brigther: B4_DC from 2.61 to 2.54 mW; B4_12 to 0.135 from 0.138 mW (see again fig.3). In the meantime, B1p Phase cameras started to get worse a lot (especially the two e UPP and LWR 6 MHz sidebands, see Fig. 6).

9.57.26 UTC: CITF unlock.


%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%


10.53.20 UTC: lock CITF -> 12 MHz 0.09 mW; B4_DC 1.95; B4 phase cameras had a very good circular shape, while B1p cameras kept getting worse (see Fig. 7). During this lock at first the BS was fine aligned by hand to improve the figures of merit, but we immediately noticed that the BS AA didn't engage as the QPDs shutters were closed, so we opened them by hand. The situation sligthly improved.

then at around 12.05 the CITF seemed to almost stabilize with constant figures of merit value, so we started to perform an other set of scans.

However, this last set of scan didn't succed, as everytime we unlocked as we were passing by the 56MHz LWR HOM2 mode, i.e. to the corresponding value of around 21500 Hz of CARM offset. We tried different ramp values, or adjusting MICH gain (as its correction sometimes saturated), but we never went past that value.

For reference the GPS of the scan attempts are the following:

12.05.45 CARM set 35.000 in 60 sec.
12.06.22 unlock CITF


12.12.16 UTC: scan 35000 Hz offset, 300 sec ramp;
12.15.10 unlock CITF at 21500 Hz offset;


12.18.00 UTC: scan 35000 Hz offset, 300 sec ramp;
12.20.57 unlock CITF at 21500 Hz offset;


12.27.59 UTC: scan 35000 Hz offset, 300 sec ramp;
12.30.54 unlock CITF at 21500 Hz offset; MICH correction saturated;

12.36.49 UTC: scan 35000 Hz offset, 300 sec ramp;  (this time with a lower MICH gain)
12.39.45 unlock CITF at 21500 Hz offset; This time MICH correction didn't saturate, but we unlocked anyway.

13.01.12 UTC: scan 35 kHz in 120 sec ramp.

13.02.27 UTC: unlock CITF at 21500 Hz offset;


%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

After these failed attempts, we relocked the CITF with the usual negative offset (-3000 Hz), in order to explore a new Chrocc power value.

At 13.39 UTC Diana started a Hartman measurement to take a reference.

At 13.47.53 UTC Chrocc Voltage has been set to a higher value: 0.4 V.

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

In the following table are reported the different Chrocc settings performed in the past shifts, in terms of Power (W), Voltage (V) and Dioptres, together with the two equations to obtain the conversion from Voltage to Power and to Power to Dioptres.

Chrocc settings:
  P [W] V_dac [V] Dioptres
Nominal 0.78 0.43 2.23e-6
August 04th tuning 0.98 0.50 2.8e-6
August 06th 1st tuning 0.61 0.36 1.74e-6
August 06th 2nd tuning 0.71 0.40 2.03e-6

Chrocc Power:

P\left ( W \right ) = 1.86Vdac\left ( V \right )^{2} + 1.022Vdac\left ( V \right ) + 9.55e^{-11}
Dioptres:

D = 2.859e^{5}\cdot P

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

At the end of the shift we performed a scan while locked on beating signals, to evaluate the mismatch percentage. GPS of the scan are:

For the TEM 00 mode: 15.49.10 UTC

For the TEM 02 mode: 15.53.27 UTC.

Results are reported in Fig. 8., where a 3 % of mismatch has been found on North Arm.

We left the ITF locked on the IR, with the Chrocc left to the last voltage setting (i.e. 0.4 V).

Images attached to this report
Comments to this report:
jerome.degallaix - 9:14 Monday 08 August 2022 (56683) Print this report

As noticed in the post, the CHRoCC voltage of 0.43V is a really good starting point. As displayed in the VIM pages, the SB power recycling gain for the 6 MHz was ~65 and ~40 for the 56 MHz. It has to be compared to 40 and 20 without the CHRoCC (see VIR-0369A-22 for example)

Images attached to this comment
Detector Operation (Operations Report)
menzione - 16:10 Saturday 06 August 2022 (56671) Print this report
Operator Report - Daily shift

Upon arrival I found ITF locked on IR. I locked CITF realigning PR and SR (cross) then, I easily locked ITF up to CARM_NULL_3F at the first attempt.
At 07:00 UTC, the planned ISC activity on – CHRoCC power tuning until CITF lock (Bersanetti, Pinto, Lumaca) started and went on without major problems performing scans with ITF locked at LOCKEDARMS_BEAT_DRMI_1F.
Activity is still in progress.

No particular operations to report.
No DMS Events to report

AdV-DET (Phase camera (2 to be developed))
beuzekom, tacca - 11:28 Saturday 06 August 2022 (56670) Print this report
B1p phase camera larger beam
Following the request of the commissioning coordinator, we increased the size of the beam acquired by the B1p phase camera. The current value for the spiral radius is 0.006. The size of the beam will be adjusted again once the interferometer will be locked in CARM NULL to avoid clipping on the carrier image.
Images attached to this report
Detector Operation (Operations Report)
menzione - 8:27 Saturday 06 August 2022 (56669) Print this report
Operator Report - Initial Shift

ITF State: ITF Locked on IR.
ITF Mode: LOCKING
Quick Summary: IMC and RFC locked; all Suspensions loops closed, all SBE loops closed.
Activities ongoing since this morning: No activity ongoing

No DMS Events to report

Sub-system reports

EnvMoni
METEO station down since yesterday at 1:30 UTC (plot attached), "Esternal, FlatChannel, Wind activity" red flags on DMS.

Images attached to this report
AdV-TCS (CO2 laser projector)
gherardini, lumaca - 0:09 Saturday 06 August 2022 (56668) Print this report
NI CH center position check with NI RH and HWS DET

The goal of the activity was to check the relative position of the NI RH center (i.e. the NI mirror center) and the NI CH center, continuing the activity started in 56633.

In preparation for the shift, at 14:22 UTC the NI DASes were blocked on the bench by flip mirrors and the ITF was realigned on green, after the work performed during the morning shift.

At 16.25 UTC Fabio set the ITF in N arm locked, W arm misaligned.

After checking the HWS DET illumination, at 16.26 UTC an HWS DET acquisition started and at 16.36 UTC the NI RH was switched ON @ 15 W. 

After almost 1 hour the NI RH induced curvature center starts to stabilize: results are shown in table 1 and fig.1. The NI RH was switched OFF at 18.14 UTC.

 

NI RH (03/08)

NI RH (04/08)

difference

X [m]

(0.0003 +- 0.005)

(0.023 +- 0.015)

23 mm

Y [m]

(-0.008 +- 0.004)

(-0.007 +- 0.002)

1 mm

The NI RH center position changed by around 23 mm along the horizontal direction: it could be justified by the intervention of the morning shift (56654), during which the position of the meniscus lens (through which also the HWS DET SLED beam passed) was changed. It could be worth considering a realignment of HWS DET SLED beam on top of NI TM.

At 20.06 UTC a new HWS DET acquisition was started and at 20.15 UTC the NI CH was switched ON at high power. The induced curvature became immediately evident and soon after the center stabilized: the comparison results between NI CH and RH centers are shown in table 2 and fig.2. At 20.27 UTC the NI CH was switched OFF.

 

NI RH (05/08)

NI CH (05/08)

difference

X [m]

(0.023 +- 0.015)

(-0.0066 +- 0.0019)

29 mm

Y [m]

(-0.007 +- 0.002)

(-0.003 +- 0.002)

4 mm

Still a slight mis-centering along horizontal direction was present. I performed an horizontal movement to recover this mis-alignement: it should be checked in a next shift.


At the end of the shift, the NI DASes actuators were switched ON again ( 21.22 UTC), at the powers left from last shift (56650) to compensate for the NI CP cold lens, and the ITF was left locked on IR.

Images attached to this report
Comments to this report:
nardecchia - 20:05 Sunday 07 August 2022 (56677) Print this report

The check of the last movement has been done today afternoon.

With the ITF in down, the NI CH has been switched ON at high power and the minimum OPL coordinates have been compared with the maximum RH OPL ones (see the attached figure).

The results are summarized in the following table:

 

X [pixels]

Y [pixels]

NI RH

729+-6

459+-4

NI CH

727+-4

464+-24

Difference

2+-10

5+-28

Thus, NI CH and RH are centered within 12*12.28/1024*11~2 mm in X and within  33*12.28/1024~4.5 mm in Y.

The ITF has been left locked on IR.

 

Images attached to this comment
nardecchia - 15:00 Monday 08 August 2022 (56686) Print this report

I have to correct the sentence:

''Thus, NI CH and RH are centered within 12*12.28/1024*11~2 mm in X and within  33*12.28/1024~4.5 mm in Y.''

in

''Thus, NI CH and RH are centered within 12*12.28/1024*13~2 mm in X and within  33*12.28/1024*13~5 mm in Y.''

I have mistakenly used the HWS-INJ magnification (11) instead of the HWS-DET one (13).

Detector Operation (Operations Report)
gherardini - 23:10 Friday 05 August 2022 (56666) Print this report
Operator Report - Afternoon shift

This afternoon shift was dedicated to NI CO2 actuator centering check, the work work was carried out with the north cavity locked and the west misaligned, it went on for the whole shift without any major problem; activity concluded at the end of the shift, we left the cavities locked on the infrared.

Sub-system reports

Other
at 15:19UTC the SQB1 position control was opened by the guardian; control closed after re adjust of F1 vertical position with stepping motors.

Vacuum
west end LN2 tank extraordinary refill from 15:30UTC to 16:45UTC

Images attached to this report
AdV-SGD (FDS commissioning)
bawaj, guo, vardaro - 22:33 Friday 05 August 2022 (56656) Print this report
SQZ_FLT: bugfix and preparation for a long test run

We successfully ported PyPllServer to the default cmt-conda environment.

After that, we prepared an automation branch (dev-mb_220805_for_long_run) adjusted to the long test run. With respect to the currently most updated dev-mb branch it contains changes:
- PySQZ_IR is enabled all the time
- LFC_UGF decorator is disabled (verify if this decorator or LOCK_FLAG creates a problem in case of lock at low GR transmission).

In the process we realized that a measurement of the IR AA sensing matrix might be necessary. We made an attempt but due to the presence of strong wind we were not successful. On the other hand we dedicated time to write down the sensing matrix measurement algorithm which you can find beneath.

The automation was started at (useful data since) 1343766188 05 Aug 2022 20:22:50 UTC

IR AA sensing matrix measurement procedure

Prerequisites:

  1. Cavity is AOM_LOCKED_IR, (IR_AA is off).
  2. Demodulation Phase of the error signals must be correct.
  3. Error signal noise below: ?
  4. IR Galvos must be ON and OK

Actuators

  • EQB2_IR_QD1_RF_11MHz_X
  • EQB2_IR_QD1_RF_11MHz_Y
  • EQB2_IR_QD2_RF_11MHz_X
  • EQB2_IR_QD2_RF_11MHz_Y

Error signals

  • EQB1_IR_M3_MT1_X
  • EQB1_IR_M3_MT1_Y
  • EQB1_IR_M5_MT1_X
  • EQB1_IR_M5_MT1_Y

Steps:

  1. Put the offset on one actuator (command is in VPM)
  2. Wait one second
  3. Read average values of all error signals
  4. Invert the offset on the actuator
  5. Wait one second
  6. Read average values of all error signals (detrend?)
  7. Calculate the differences between the two readouts and divide by the difference in offsets.
  8. Put results to the matrix column.
  9. Repeat steps from 1-8 for other three actuators.
  10. Invert the measured matrix
Comments to this report:
vardaro - 0:28 Monday 08 August 2022 (56679) Print this report

The automation worked well up to 14:48 UTC of 7th August, then the filter cavity unlocked and started to lock/unlock up to 15 UTC see Fig1. A detailed entry to this behaviour will follow/

After that we experienced an INJ unlock and the ITF ML set point moved by about 0.3V (see Fig2). After that the PLLs were no more able to lock thus I had to tune the SQZ laser set point to lock the PLL in this working point (new set points in Fig3). Then I realigned the Green beam and the IR beam to the filter cavity and I re engaged the automation at 21:45 UTC see Fig4.

Images attached to this comment
vardaro, bawaj - 17:22 Monday 08 August 2022 (56689) Print this report

After the restore of yesterday the Automation worked well up to 6:15 UTC of this morning. Here the cavity unlocked and was not more able to relock again (fig1) . We did the following observation:

  • after the unlock the system fall in down and discharged the correction of all the mirrirs (FCIM, FCEM GR_M5 and GR_M7) but some kicks to the mirror occured. After that the transmission was at 4V maximum (fig2)
  • the threshold for the lock was set at 6.2 V thus the system did not relock on green anymore
  • sometimes after the lock on green we noticed a lot of frequency noise and the BPC2 based on dither lines on FCIM and FCEM did not work well. 

To solve it we are implementing the following solutions:

  • increase the SQZ_CC_Ctrl GAIN from 0.12 to 0.5 to reduce frequency noise (temorary bug fixing)
  • lower the threshold for engage the lock on green from 6.2 to 5 i.e. the 75% of the maximum transmission
  • insert in the green+IR locking a state to re-align the green if the transmission is too low, because in that condition it is difficult to find the IR
  • try to discharge the mirrors correction without sending kicks to the actuators. A possible solution is tested (fig 3).

We are impementing these upgrades in the SQZ_FLT.py code and tomorrow we will test them. 

Note that other unlocks were triggered by other issues in the automation. These kind of unlocks was less critical thus we started from the bigger problem

Images attached to this comment
AdV-DET (Commissioning)
gouaty, hui, mwas, sposito - 19:19 Friday 05 August 2022 (56654) Print this report
OMC mode matching and benches alignment checks

This morning we worked on the OMC mode matching tuning, and performed some alignment checks on SPRB and SIB2.

First Gianmatteo put the ITF in NI single bounce in order to perform the activities on SPRB and OMC mode matching.

We started by recentering the B4 beam on the camera (it was slightly off centered) at the beginning of the shift. SPRB new angular angular setpoints: TX=156, TY=-663

We opened the OMC shutter using DET_MAIN, then we tested the shutter closing and we saw that it was not working: Michal found a bug in the graph of the node, as two paths are possible to go from "SHUTTER CLOSED" to "SHUTTER OPEN". The problem was linked to the fact that metatron was chosing a different path than the one it used to. Michal fixed this issue by increasing the length of the path that we do not want metatron to select. Then the OMC shutter closing worked successfully.

We performed a first scan of the OMC at 07h08m40 utc to check the initial status of alignment: TEM00 peak 0.04mW, order 1 peak at 0.02 mW. Thus the OMC was strongly misaligned.
OMC locked at 07h41m20 utc, and realigned using the picomotors on SDB1.

Another OMC scan at 07h55m40 utc to check the initial status of mode mismatch: TEM00 0.050 mW, order 1: 0.0013 mW (mostly horizontal) > 2.6% misalignment defect, order 2: 0.008 mW > 16% mode mismatch.

We started the mode matching tuning by moving the SDB1 bench along the longitudinal axis by acting on the suspension top stage (F0_X). Our starting point corresponds to: Sa_OB_F0_X = -2200 and SDB1_LC_Z = -64 um. Bringing the suspension top stage at 0 (F0_X = 0) we noticed that the mode matching was improving (the transmission through the OMC was increasing as well as the normalized signal B1_PD3_norm_B1s).

Then we moved the meniscus lens SDB1_L1_Z in the negative direction (which corresponds to positive direction for the bench). In parallel we reajusted the balancing of the bench by acting on the motorized counter-weight (BENCH_TX). Below are the details of the steps we performed:

SDB1_L1_Z -60000 -80000 -70000 -80000 TOTAL: -290000 steps
BENCH_TX (counter weight) -150000 -140000 -130000 -140000 TOTAL: -560000 steps

We realigned the OMC after the first -140000 steps and after the -290000 steps. Each time we noticed a clear improvement of the OMC mode matching.

Performing -60000 steps with the lens seemed to be equivalent to +500 um with the suspension for the mode matching. Thus we estimated that the meniscus lens was displaced by ~2.5 mm. Concerning the motorized counter-weight, it is likely that we have used most of its available range, thus we preferred to stop moving the lens farther to not risk blocking the counter-weight if we reach the end of range.

We continued improving the mode matching by moving the suspension top stage towards positive X values. Fig.1 shows the overall trend of the OMC transmission during the tuning of the mode matching (the meniscus lens was moved between 8h30 and 9h30 utc).

Sa_OB_F0_X at +2500:
carrier TEM00 reaching about 0.062 mW.
OMC scan (with rising up temperature) started at 10h05m20 utc:
Order 1: small
Order 2: 0.0011 mW > 1.8% mode mismatch
New scan (with decreasing temperature) at 10h14m16 utc:
order 1: 0.00025 mW > 0.4% of misalignment defect
order 2: 0.0012 mW > 1.9% mode mismatch

Sa_OB_F0_X at +4500:
Scan of the OMC at 11h14m00 utc:
order 1: 0.00042 mW > 0.7% misalignment
order 2: 0.0003 mW > 0.5% mode mismatch
Scan decreasing at 11h21m30 utc (integration time of B1 camera increased at maximum) - see Fig.2:
order 2: 0.0003 mW (the mode has a nice shape of 4 leaf clover - see Fig.3).
order 1: 0.00046 mW

As the suspension was close to its end of range (+5000), and the bench LVDT were quite off-centered (SDB1_LC_Z = 5500um), this bench position did not seem optimal for the local control, thus we decided to come back to the suspension position Sa_OB_F0_X at +2600 um. We keep this as nominal setpoint for the time being.

After realigning the OMC we performed a final scan at 11h55m00 utc (see Figure 4):
Order 1: 0.0007 mW
Order 2: 0.0009 mW > 1.5% mode mismatch

Conclusions on the OMC mode matching tuning: We demonstrated that a mode mismatch of 0.5% can be reached. However, in order to limit the off-centering of the suspension and bench LVDT we decided to stay at a working point where the mode mismatch should be at the level of 1.5-2%. In the future we would like to move further the menicus lens is that is possible.

After the OMC mode matching activity, we closed the OMC shutter and updated the SDB1 angular setpoint in the config file.

 

Then Gianmatteo aligned the PR mirror and we could check the alignment of the photodiodes on SIB2 bench:
To this purpose we scanned the SIB2 bench in TY: Fig.5 and in TX: Fig.6.

For the current bench angular setpoints (TX = -100 and TY = -325 urad) the beam seems well centered inside the B2 photodiodes (we are close to the center of the plateau visible in the TY, and the TX scan does not change the power on these photodiodes). For the RFC photodiodes, we observe a slow trend as we displace the bench. It does not look like the beam going out from the photodiode (no sudden drop). I wonder if it could be related to the beam size and shape (mode of order 1), which could be slightly clipped, or some junk light (reflection on AR coating for PD2), or maybe a polarization effect.

We checked that the galvo loops of B2_QD2 could be closed with the DC centering. Switched back to 2f centering after the test.

 

At the end of the shift, we updated the current limit threshold of the B4 photodiodes to 90 instead of 80 mA, to try to solve the issue mentioned yesterday: https://logbook.virgo-gw.eu/virgo/?r=56646 . SPRB configuration reloaded at 14h05m38 utc.

Images attached to this report
Comments to this report:
gouaty, hui, mwas - 9:55 Monday 08 August 2022 (56684) Print this report

During the activity of friday morning (OMC mode matching), the bench angular setpoints have evolved from 136 to 29 urad in TX , and from 25 to 32 urad in TY. The change in TY seems rather unsignificant, but the change in TX is quite substantial. We see that at the same time the beam spot is moving on the B1p camera. The main effect is a horizontal motion on the camera by 300 um.

Considering the size of the photodiodes (3 mm aperture), a motion of 300 um should not be an issue, but it may still be worth to check the centering of the beam on the photodiodes and the quadrants on SDB2.

Also probably the alignment of the B1s QPD on EDB will have to be checked.

Images attached to this comment
Search Help
×

Warning

×